# GroundingDINO-US-SAM: Text-Prompted Multi-Organ Segmentation in Ultrasound with LoRA-Tuned Vision–Language Models

### v2

@@ 3.3 Generalization to Unseen Datasets
As shown in Table 4, our method achieves the highest average DSC ($72.83\%$) and IoU ($68.53\%$) across datasets, outperforming all competing approaches without retraining or domain-specific adaptation. On the BUSBRA dataset, our model reached a DSC of $79.10\%$ and IoU of $86.44\%$, exceeding all other methods. For the TNSCUI thyroid dataset, it achieved the highest DSC ($86.44\%$) and IoU ($79.10\%$), showing strong robustness. On the Luminous dataset, while SAMUS slightly outperformed our method in DSC (65.36% vs 64.99%), our approach achieved the best IoU ($51.96\%$ vs $49.99\%$), indicating better overlap with the ground-truth masks.
Finally, it is worth noting that SAMUS and BiomedParse benefit from much larger pretraining datasets (30,000 and 3.4 million images, respectively), whereas our model is trained on significantly less data yet generalizes more effectively across diverse anatomical structures and imaging conditions.
Table 4: Quantitative comparison of segmentation performance (DSC and IoU scores) on unseen public ultrasound datasets across different organs using SOTA methods: UniverSeg, BiomedParse, SAMUS, MedCLIP-SAM, MedCLIP-SAMv2, and the proposed method. Results are presented as mean $\pm$ standard deviation. Bolded results indicate the highest score for each comparison.
@@
Breast BUSBRA 47.97±27 35.72±24 78.79±22 69.05±23 75.04±15 64.80±17 53.58±25 40.47±23 57.82±38 49.93±35 79.10±18 86.44±18
Thyroid TNSCUI 6.21±7 3.37±4 40.54±39 33.62±34 63.70±23 57.50±24 63.90±13 48.17±13 76.34±12 63.06±14 86.44±17 79.10±20
Multifidus Muscle Luminous 25.54±16 15.69±11 12.59±18 7.86±12 65.36±13 49.99±15 52.27±25 38.96±22 62.96±24 49.47±21 64.99±24 51.96±22
Average 26.57±16 18.26±13 43.97±26 36.84±23 68.03±17 57.43±18 56.58±21 42.53±19 65.71±24 54.15±23 72.83±20 68.53±20
@@ 3.4 Results with Different Text Prompts
Table 5 shows how segmentation performance varies with different text prompts for each kidney structure. Results indicate that the model is generally robust to prompt wording, with minor performance fluctuations. For example, for the Cortex, the DSC values varied between 70% and 72% for five different prompts. Prompts containing terms widely recognized in clinical contexts (“renal capsule”, “renal cortex”) consistently perform well. Also, the capsule segmentation achieves higher performance (DSC 85–92%, IoU 77–85%) compared to the cortex, medulla, and central echo complex (DSC 66–73%, IoU 50–58%), suggesting that well-defined, clear anatomical boundaries (such as the capsule) are easier for SAM2 to segment accurately.
Table 5: Segmentation performance (DSC and IoU scores) of different kidney structures using various text prompts. Results are reported as mean $\pm$ standard deviation.
Structure Prompt Text DSC (%) $\uparrow$ IoU (%) $\uparrow$
Cortex segment kidney cortex 70.01±14 55.46±15
delineate cortex 72.04±11 57.32±12
segment renal cortex 71.95±10 57.14±12
cortical region of the kidney 72.55±8 57.48±9
@@
delineate renal capsule 91.71±5 85.05±8
outline kidney capsule boundary 89.72±9 82.31±12
capsular layer of the kidney 85.44±14 76.93±19
identify renal encapsulation 91.17±5 84.18±8
Medulla segment kidney medulla 69.85±17 55.77±17
delineate renal medullary region 69.01±17 54.91±17
outline medulla of the kidney 69.69±12 54.69±13
medullary area of kidney 68.82±15 54.35±17
identify renal medulla 70.32±16 56.20±16
@@
delineate kidney central echo complex 68.45±14 53.77±16
renal central echo complex 67.23±15 52.50±16
outline central echoes of kidney 65.67±14 50.52±16
delineate renal central echo complex 67.56±16 52.97±17
@@ 3.6 Runtimes
Table 6: Average inference time per image over 10 runs on NVIDIA TITAN V with 12 GB of RAM.
@@
UniverSeg 0.21
BiomedParse 0.49
SAMUS 0.67
MedCLIP-SAM 3.05
MedCLIP-SAMv2 2.73
Ours 0.33
@@ 4 Discussion
BiomedParse and SAMUS benefited from significantly larger pre-training datasets, yet our model, fine-tuned on a much smaller dataset, consistently achieves the best segmentation results. It should also be noted that SAMUS requires a point prompt at test time, and in our evaluation, we supplied a randomly chosen point inside each ground-truth mask. Our method, in contrast, operates using solely text and receives no spatial guidance, making the evaluation deliberately conservative and underlining the method’s robustness. These findings provide two insights: (1) A Grounding DINO encoder pre-trained almost entirely on natural-image/text pairs can be tuned to the ultrasound domain with minimal data. In other words, LoRA fine-tuning is enough to align visual and language cues. (2) The SAM2 decoder is capable of high-quality ultrasound segmentation, but only when an accurate spatial prompt is available; without such guidance, its advantage disappears (as shown by the poor results of Grounding DINO + SAM2 in Table 2).
@@ 6 Acknowledgment
Funded by Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery grant. We would like to thank NVIDIA for the donation of the GPU.

### v3

@@ 3.3 Generalization to Unseen Datasets
As shown in Table 5, our method achieves the best performance across most benchmarks. On the BUSBRA breast dataset, it achieves the top DSC of 79.10% and IoU of 86.44%. On the TNSCUI thyroid dataset, our model again outperforms all others, reaching a DSC of $86.44\%$ and IoU of $79.10\%$. Although SAMUS slightly outperforms our method in DSC on the Luminous dataset ($65.36\%$ vs. $64.99\%$), our model achieves the highest IoU of 51.96%, indicating stronger alignment with the ground truth masks. Notably, both SAMUS and BiomedParse benefit from large-scale training data—30,000 and 3.4 million images, respectively. In contrast, our model is trained on significantly fewer samples and still demonstrates strong generalization across different anatomical structures and imaging conditions, without retraining or task-specific adaptation.
Table 5: Quantitative comparison of segmentation performance (DSC and IoU scores) on unseen public ultrasound datasets across different organs using SOTA methods: UniverSeg, BiomedParse, SAMUS, MedCLIP-SAM, MedCLIP-SAMv2, and the proposed method. Results are presented as mean $\pm$ standard deviation. Bolded results indicate the highest score for each comparison. Statistical significance was assessed using paired t-tests: ∗∗ denotes $p<0.005$ and denotes $p<0.01$.
@@
Breast BUSBRA $47.97±27^{**}$ $35.72±24^{**}$ $78.79±22^{**}$ $69.05±23^{**}$ $75.04±15^{**}$ $64.80±17^{**}$ $53.58±25^{**}$ $40.47±23^{**}$ $57.82±38^{**}$ $49.93±35^{**}$ 79.10±18 86.44±18
Thyroid TNSCUI $6.21±7^{**}$ $3.37±4^{**}$ $40.54±39^{**}$ $33.62±34^{**}$ $63.70±23$ $57.50±24$ $63.90±13^{**}$ $48.17±13^{**}$ $76.34±12$ $63.06±14$ 86.44±17 79.10±20
Multifidus Muscle Luminous $25.54±16^{**}$ $15.69±11^{**}$ $12.59±18^{**}$ $7.86±12^{**}$ 65.36±13 $49.99±15$ $52.27±25^{*}$ $38.96±22^{**}$ $62.96±24$ $49.47±21$ $64.99±24$ 51.96±22
@@ 3.4 Results with Different Text Prompts
Table 6 presents the segmentation performance (DSC and IoU) for various anatomical structures in the kidney and thyroid, using a range of text prompts. The results demonstrate that prompt phrasing has a considerable impact on segmentation accuracy. Well-designed prompts that are anatomically accurate and concise, such as “segment kidney cortex” or “segment the thyroid nodules”, achieved high Dice Similarity Coefficient (DSC) and Intersection over Union (IoU) scores across both organs. For instance, segmenting the kidney capsule with the prompt “segment kidney capsule” yielded a DSC of 91.74% and IoU of 85.10%, while the thyroid parenchyma using “segment the glandular portion of the thyroid” achieved a DSC of 77.07% and IoU of 64.73%. In contrast, prompts that were ambiguous, anatomically irrelevant, or unrelated to the target structure (highlighted in gray in the table) resulted in significantly lower performance. Examples include prompts such as “highlight the middle phalanges” or “segment zones of the uterus”, which consistently yielded suboptimal scores regardless of the target organ. This drop in performance highlights the model’s reliance on the semantic alignment between the prompt and the visual content. Overall, these results highlight the crucial role of a precise and contextually relevant prompt formulation. To ensure consistency and clinical relevance, we consulted with a radiologist when designing the primary prompt templates used in this study.
Table 6: Segmentation performance (DSC and IoU) for various anatomical structures using different text prompts. Poorly performing prompts are highlighted in gray to illustrate the limitations of ill-conceived or ambiguous phrasing. Results are reported as mean $\pm$ standard deviation.
Organ Structure Prompt Text DSC (%) $\uparrow$ IoU (%) $\uparrow$
Kidney (KidneyUS) Cortex segment kidney cortex 70.01±14 55.46±15
segment renal cortex 68.17±18 54.13±17
@@
delineate cortex 68.25±19 54.30±17
cortical region of the kidney 17.82±31 14.12±25
highlight the middle phalanges 44.35±33 34.13±26
segment zones of the uterus 41.90±37 33.78±30
@@
segment renal capsule 91.69±5 85.03±7
outline kidney capsule boundary 89.72±8 82.31±12
identify capsule 90.69±8 83.77±11
capsular layer of the kidney 41.73±43 37.57±40
highlight the middle phalanges 43.55±33 33.80±32
segment zones of the uterus 42.86±42 37.96±39
Medulla segment kidney medulla 68.79±18 68.79±18
segment renal medulla 67.83±20 54.18±18
identify renal medullary area 41.13±36 32.80±30
outline medulla 66.18±21 52.68±20
medullary area of kidney 22.94±33 18.12±27
highlight the middle phalanges 41.97±35 33.18±29
segment zones of the uterus 40.24±37 32.52±31
@@
segment renal central echo complex 68.45±14 53.77±16
delineate central echoes of kidney 64.25±21 50.46±20
outline central echoes 63.59±21 49.80±20
renal central echoes 65.71±18 51.39±18
highlight the middle phalanges 49.14±31 37.89±26
segment zones of the uterus 42.14±35 33.05±28
Thyroid (TN3K, TG3K) Parenchyma segment the thyroid parenchyma 70.32±26 59.22±24
segment the glandular portion of the thyroid 77.07±15 64.73±16
outline the thyroid tissue 74.29±20 62.37±20
delineate parenchyma 31.64±38 26.36±33
glandular area of the thyroid 75.32±19 63.39±19
highlight the middle phalanges 55.11±37 46.34±32
segment zones of the uterus 50.88±25 49.46±23
Nodule segment the thyroid nodules 75.90±21 65.02±22
segment the focal lesions in the thyroid 75.10±21 64.06±22
identify the thyroid lump 72.13±26 61.57±25
outline nodules 72.04±27 61.83±26
hypoechoic region in the thyroid 33.59±41 29.54±37
highlight the middle phalanges 51.96±30 45.21±28
segment zones of the uterus 45.94±33 43.78±30
@@ 3.6 Runtimes
Table 7: Average inference time per image (mean ± standard deviation) over 10 runs on NVIDIA TITAN V with 12 GB of RAM.
@@
UniverSeg 0.21±0.002
BiomedParse 0.49±0.045
SAMUS 0.67±0.005
MedCLIP-SAM 3.05±0.099
MedCLIP-SAMv2 2.73±0.085
Ours 0.33±0.032
@@ 4 Discussion
BiomedParse and SAMUS benefited from significantly larger pre-training datasets, yet our model, fine-tuned on a much smaller dataset, consistently achieves the best segmentation results. It should also be noted that SAMUS requires a point prompt at test time, and in our evaluation, we supplied a randomly chosen point inside each ground-truth mask. Our method, in contrast, operates using solely text and receives no spatial guidance, making the evaluation deliberately conservative and underlining the method’s robustness. These findings show that Grounding DINO encoder pre-trained almost entirely on natural-image/text pairs can be tuned to the ultrasound domain with minimal data. In other words, LoRA fine-tuning is enough to align visual and language cues. When comparing the poor results of Grounding DINO + SAM2 to the excellent results of Fine-tuned Grounding DINO + SAM2 in Table 3, it is evident that the SAM2 decoder is capable of high-quality ultrasound segmentation when an accurate bounding box is available. In other words, this implies that fine-tuning is required solely for Grounding DINO, with no modifications needed for SAM2. While our preliminary results show that it is not necessary to fine-tune SAM2 to the ultrasound domain, we will explore this fine-tuning in future work.
@@ 6 Acknowledgment
Funded by Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery and Government of Canada’s New Frontiers in Research Fund (NFRF) [NFRFE-2022-00295] grants. We would like to thank Dr. Khashayar Rafatzand for invaluable discussions and NVIDIA for the donation of the GPU.